# Inverse thermodynamic uncertainty relation and entropy production

### v1

@@ Abstract
where $\tilde{\lambda}:=\sqrt{\lambda(\lambda+2\kappa)}$. Note that the iTURs for the discrete variable case cannot reproduce the iTURs for continuous variable case since the escape rate $\kappa$ diverges in the continuous variable limit. In the SM [59], we provide the proof for a tighter iTUR [61].
Outline of the proof.— We here outline the proof, while the details are presented in the SM [59]. Utilizing a variational principle for diffusion coefficients, we can express $D_{\infty}$ as
$D_{\infty}$ $=\sup_{\bm{\chi}}\inf_{\bm{\eta}}\bigg{[}2\,\text{Cov}_{\text{st}}\left((\bm{\theta}+\nabla\bm{\eta})\cdot\bm{\nu}_{\text{st}},\bm{\chi}\right)$ (15)
$\quad+\langle(\bm{\theta}+\nabla\bm{\eta})\cdot\bm{B}(\bm{\theta}+\nabla\bm{\eta})\rangle_{\text{st}}-\langle\nabla\bm{\chi}\cdot\bm{B}\nabla\bm{\chi}\rangle_{\text{st}}\bigg{]},$
where $\langle\diamond\rangle_{\text{st}}$ and $\text{Cov}_{\text{st}}(\diamond,\diamond)$ denote the steady-state average and covariance, respectively. Setting $\bm{\eta}(\bm{x})=\bm{0}$ simplifies Eq. (15) to
$D_{\infty}$ $\leq D_{0}+\sup_{\bm{\chi}}\left[2\,\text{Cov}_{\text{st}}(\bm{\theta}\cdot\bm{\nu}_{\text{st}},\bm{\chi})-\langle\nabla\bm{\chi}\cdot\bm{B}\nabla\bm{\chi}\rangle_{\text{st}}\right],$ (16)
where $D_{0}=\langle\bm{\theta}\cdot\bm{B}\bm{\theta}\rangle_{\text{st}}$. Using the Cauchy–Schwarz inequality, we obtain
$D_{\infty}\leq D_{0}+\sup_{\bm{\chi}}\left[\frac{\left[\text{Cov}_{\text{st}}(\bm{\theta}\cdot\bm{\nu}_{\text{st}},\bm{\chi})\right]^{2}}{\langle\nabla\bm{\chi}\cdot\bm{B}\nabla\bm{\chi}\rangle_{\text{st}}}\right].$ (17)
We define the symmetrized Fokker–Planck operator, $\mathbb{L}_{\rm sym}$, by substituting the original drift in $\mathbb{L}$ with the modified field $\tilde{\bm{A}}(\bm{x})=\bm{B}\nabla\ln p_{\text{st}}(\bm{x})$. This substitution renders $\mathbb{L}_{\rm sym}$ symmetric with respect to a particular inner product weighted by the steady-state distribution. Its spectral gap, $\lambda$, can then be determined via the variational formula:
$\lambda=\inf_{\chi}\left[\frac{\langle\nabla\chi\cdot\bm{B}\nabla\chi\rangle_{\text{\rm st}}}{\text{Var}_{\text{\rm st}}(\chi)}\right].$ (18)
We can deduce that $\lambda$ is upper bounded by the ratio $\langle\nabla\chi\cdot\bm{B}\nabla\chi\rangle_{\text{st}}/\text{Var}_{\text{st}}(\chi)$. This leads to the bound:
$D_{\infty}\leq D_{0}+\frac{1}{\lambda}\text{Var}_{\text{st}}(\bm{\theta}\cdot\bm{\nu}_{\text{st}})\,.$ (19)
We now focus on bounding $\text{Var}_{\text{st}}(\bm{\theta}\cdot\bm{\nu}_{\text{st}})$. By definition, we have
$\text{Var}_{\text{\rm st}}(\bm{\theta}\cdot\bm{\nu}_{\text{\rm st}})=\langle(\bm{\theta}\cdot\bm{\nu}_{\text{\rm st}})^{2}\rangle_{\text{\rm st}}-j^{2}_{\text{\rm st}}\,.$ (20)
The first term can be bounded as follows:
$\langle(\bm{\theta}\cdot\bm{\nu}_{\text{\rm st}})^{2}\rangle_{\text{\rm st}}\leq\langle(\bm{\theta}\cdot\bm{B}\bm{\theta})(\bm{\nu}_{\text{\rm st}}\cdot\bm{B}^{-1}\bm{\nu}_{\text{\rm st}})\rangle_{\text{\rm st}}\,.$ (21)
Introducing $D_{\text{max}}$, satisfying $\bm{\theta}\cdot\bm{B}\bm{\theta}\leq D_{\text{max}}$, and noting $\sigma_{\text{st}}=\langle\bm{\nu}_{\text{st}}\cdot\bm{B}^{-1}\bm{\nu}_{\text{st}}\rangle_{\text{st}}$, we obtain
$\text{Var}_{\text{st}}(\bm{\theta}\cdot\bm{\nu}_{\text{st}})\leq D_{\text{max}}\sigma_{\text{st}}-j_{\text{st}}^{2}\,.$ (22)
Substituting this bound into Eq. (19), we arrive at the iTUR bound (10) with $\tau\to\infty$.
@@
Summary and Discussion.— We have derived upper bounds on current fluctuations at steady state (iTUR) by utilizing both the entropy production rate and the spectral gap. By integrating the thermodynamic uncertainty relation (TUR), our results constrain the range of current fluctuations. While the TUR suggests that current fluctuations relative to the average current can be reduced by increasing the entropy production rate, our iTUR involving entropy production implies that the overall extent of fluctuations tends to increase. Crucially, this effect is not apparent in the iTUR derived in Ref. [44]. We apply the derived iTUR to the phenomenon of giant diffusion, demonstrating the critical roles of both the spectral gap and the entropy production rate.
The iTURs introduced in this paper have significant potential for broad applications across relevant systems. For instance, the iTUR provides thermodynamic concentration inequalities, imposing upper limits on the probability of observing current deviations significantly far from their mean. Applications include not only enhancing our understanding of the underlying mechanisms, but also facilitating the estimation of maximum fluctuations and the design of efficient transport performance.
@@ Acknowledgments
This work is supported by JSPS KAKENHI Grant No. JP23K25796. AD is supported by JSPS KAKENHI (Grant No. 22K13974 and 24H00833).

### v3

@@ Abstract
where $D_{0}:=\sum_{k,m}\pi_{m}W_{mk}d_{mk}^{2}/2$ and $\tilde{\lambda}:=\sqrt{\lambda(\lambda+2\kappa)}$. Note that the iTURs for the discrete variable case cannot reproduce the iTURs for the continuous variable case since the escape rate $\kappa$ diverges in the continuous variable limit. In Appendix B of the end matter and SM [60], we provide a refined iTUR and the proof.
@@
Summary and Discussion.— We have derived upper bounds on current fluctuations at the steady state (iTUR) by utilizing both the entropy production rate and the spectral gap. By integrating the thermodynamic uncertainty relation (TUR), our results constrain the range of current fluctuations. While the TUR suggests that current fluctuations relative to the average current can be reduced by increasing the entropy production rate, our iTUR involving entropy production implies that the overall extent of fluctuations tends to increase. Crucially, this effect is not apparent in the earlier version of iTUR derived in Ref. [44]. We apply the derived iTUR to the phenomenon of giant diffusion, demonstrating the critical roles of both the spectral gap and the entropy production rate.
Although the iTUR provides a concise upper bound on fluctuations, direct experimental validation is challenging because it depends on the global suprema of local current fluctuations and on the symmetrized spectral gap—quantities that typically require detailed knowledge of the system. These challenges can be addressed by practical estimation strategies using trajectory data: supremum terms can be replaced with moment-based or probabilistic bounds, and the symmetrized spectral gap can be estimated variationally via a generalized eigenvalue procedure (see SM [60]). These strategies enable targeted tests in controlled experimental platforms such as colloidal particles in optical potentials, microfluidic devices, and single-molecule motors.
The iTURs introduced in this paper have significant potential for broad applications across relevant systems. For instance, the iTUR provides thermodynamic concentration inequalities, imposing upper limits on the probability of observing current deviations significantly far from their mean. Furthermore, the iTUR can be inverted to establish a new lower bound on the entropy production rate. This bound is complementary to the conventional TUR, proving particularly powerful for thermodynamic inference in the regime of enhanced diffusion, where the standard TUR is often less informative. The iTUR may also be useful to discuss phase transitions where current fluctuations diverge. It highlights the interplay between entropy production and the spectral gap near the transition point. Applications extend beyond theoretical insights, enabling the estimation of maximum fluctuations and providing quantitative design limits for fluctuation enhancement versus thermodynamic cost in nanotechnology and synthetic biology. Future work includes extending the iTUR beyond steady states by defining a suitable analogue of the spectral gap for time-dependent dynamics (e.g., instantaneous spectra or Floquet exponents) and formulating a quantum iTUR for open quantum systems, incorporating coherence and entropy production.
@@ Acknowledgments
The authors thank T. Van Vu and J. P. Garrahan for the insightful discussion. This work is supported by JSPS KAKENHI Grant No. JP23K25796. AD is supported by JSPS KAKENHI (Grant No. 22K13974 and 24H00833).